How many liters of a 15% sugar solution should be added to 30 liters of a 25% sugar solution
to obtain a 20% sugar solution? Solve by setting up appropriate equation(s).

Answer by Earlsdon(6294)

Let x = the required number of liters of 15% sugar solution.
The amount of sugar in this solution is 0.15x liters and the amount of sugar in the 30 liters of 25% sugar solution is 0.25(30) liters. When added together you will get a solution containing (30+x)(0.2) liters of sugar. Here's the equation you need:
0.15x+0.25(30) = 0.20(30+x) Solve for x.
0.15x+7.5 = 6+0.2x Subtract 0.15x from both sides.
7.5 = 6+0.05x Subtract 6 from both sides.
1.5 = 0.05x Divide both sides by 0.05
30 = x
30 liters of 15% sugar solution should be added to the 30 liters of 25% sugar solution to obtain 60 liters of 20% sugar solution.